OpenStudy (anonymous):

last one. What is the sum of the arithmetic sequence 140, 134, 128..., if there are 33 terms? 1,364 1,408 1,452 1,496

OpenStudy (anonymous):

Okay, we know d = 134 - 140 = -6 Applying the same reasoning as before, the sum of all the terms will be \[33a_1 + {{32(33)} \over 2}d = 33*140 + 528*(-6) = 1452\]

OpenStudy (anonymous):

thanks for all your help!!!1